# Blind-Match: Efficient Homomorphic Encryption-Based 1:N Matching for Privacy-Preserving Biometric Identification

### v1

@@ 4. Algorithm Analysis
In this section, we find the optimal $N_{in}$ for minimizing the total matching time of Blind-Match. The time taken by the Input Ciphertext Expansion algorithm (Algorithm 1), HE-$C^{3}$ (Algorithm 2), and compression method are described in Propositions 4.1, 4.2, and 4.3, respectively. Using these propositions, we will demonstrate that the total matching time of Blind-Match is minimized when $N_{in}=4$ with $m=128$, where $T_{OP,l}$ denotes the time to perform an OP (e.g., $Add$, $Mul$, $Res$, and $Rot$) operation at level $l$.
@@
The time it takes to perform Algorithm 1 can be calculated as $N_{in}\cdot(T_{Mul(P),l}+T_{Res,l})+\log(N_{in})\cdot(T_{Res,l-1}+T_{Mul(C),l-1})$.
@@
The time it takes to perform Algorithm 2 for $R$ feature vectors is $\lceil(m^{\prime}/N_{in})\rceil\cdot(N_{in}\cdot(T_{Mul(C),l}+T_{Res,l}+T_{Add,l-1})+(\log{m^{\prime}}-\log{N_{in}})\cdot(T_{Rot,l-1}+T_{Add,l-1}))$, where $m^{\prime}=(m\cdot R)/N$.
@@
The time it takes to perform the Compression method is $\lceil(m^{\prime}/N_{in})\rceil\cdot(\log{m^{\prime}}-\log{N_{in}})\cdot(T_{Mul,l}+T_{Rot,l}+T_{Add,l})$, where $m^{\prime}=(m\cdot R)/N$.
@@
Evaluating the formula for different values of $N_{in}$ with Table 1, we get: $F(2)=664.7,F(4)=434.84,F(8)=438.64,F(16)=675.6$, and $F(32)=1320.96$. Therefore, the minimum matching time is obtained when $N_{in}=4$, with a value of $F(N_{in})=434.84$.

### v2

@@ 4. Algorithm Analysis
In this section, we find the optimal $N_{in}$ for minimizing the total matching time of Blind-Match. The time taken by the Input Ciphertext Expansion algorithm (Algorithm 1), HE-$C^{3}$ (Algorithm 2), and compression method are described in Propositions 4.1, 4.2, and 4.3, respectively. Using these propositions, we will demonstrate that the total matching time of Blind-Match is minimized when $N_{in}=4$ with $m=128$, where $T_{OP,l}$ denotes the time to perform an OP (e.g., $Add$, $Mul$, $Rot$, and $Res$) operation at level $l$.
@@
The time it takes to perform Algorithm 1 can be calculated as $N_{in}\cdot((T_{Mul(P),l}+T_{Res,l})+\log(N_{in})\cdot(T_{Rot,l-1}+T_{Add,l-1}))$.
@@
The time it takes to perform Algorithm 2 for $R$ feature vectors is $\lceil(m^{\prime}/N_{in})\rceil\cdot(N_{in}\cdot(T_{Mul(C),l}+T_{Res,l}+T_{Add,l-1})+(\log{m}-\log{N_{in}})\cdot(T_{Rot,l-1}+T_{Add,l-1}))$, where $m^{\prime}=(m\cdot R)/N$.
@@
The time it takes to perform the Compression method is $\lceil(m^{\prime}/N_{in})\rceil\cdot(T_{Mul,l}+T_{Rot,l}+T_{Add,l})$, where $m^{\prime}=(m\cdot R)/N$.
@@
Evaluating the formula for different values of $N_{in}$ with Table 1, we get: $F(2)=578.02,F(4)=416.44,F(8)=429.04,F(16)=637.84$, and $F(32)=1206.04$. Therefore, the minimum matching time is obtained when $N_{in}=4$, with a value of $F(N_{in})=416.44$.